random231 (random231):

if func f(x) satisfies the relation f(x+y)=y|x-1|/(x-1)*f(x) + f(y) with f(1)=2 then

random231 (random231):

find

random231 (random231):

\[\lim_{x \rightarrow 1}f'(x)\]

random231 (random231):

@LastDayWork any idea??

random231 (random231):

@ganeshie8

OpenStudy (lastdaywork):

My answer - Zero Solution \[f(x+y) = y \frac{ |x-1| }{ x-1 }f(x) + f(y)\] substituting x = 0 ; y = 1 f(1) = -f(0) + f(1) implies f(0) = 0 \[f(x+y) = y \frac{ |x-1| }{ x-1 }f(x) + f(y)\] substituting y = 0 f(x) = f(0) = 0 Hence, f(x) = 0 ; for x ≠ 1 f(x) = 2 ; for x = 1